You wish to test the following claim (HaHa) at a significance level of α=0.002α=0.002. For the context of this problem, μd=μ2−μ1μd=μ2-μ1 where the first data set represents a pre-test and the second data set represents a post-test.

      Ho:μd=0Ho:μd=0
      Ha:μd≠0Ha:μd≠0

You believe the population of difference scores is normally distributed, but you do not know the standard deviation. You obtain pre-test and post-test samples for n=13n=13 subjects. The average difference (post - pre) is ¯d=5d¯=5 with a standard deviation of the differences of sd=25.5sd=25.5.

What is the test statistic for this sample? (Report answer accurate to three decimal places.)
test statistic = _________

What is the p-value for this sample? (Report answer accurate to four decimal places.)
p-value = _________________
Answer #1

This is a two tailed test

Ho:μd=0
Ha:μd≠0

test statistics:

dbar = 5 , sd = 25.5 , n =13

t = dbar/(sd/sqrt(n))
= 5 /(25.5/sqrt(13))
= 0.707

P value = 2 *(1-P(t < 0.707) ) with df = 12

P value = 0.4931